OpenStudy (anonymous):

What statement makes the open sentence 7x – 7 = 2x + 8 true? A. x = 0 B. x = 1 C. x = 3 D. x = 4

OpenStudy (anonymous):

7*0 – 7 = 2*0 + 8= 7*1 – 7 = 2*1 + 8= 7*3 – 7 = 2*3 + 8= 7*4 – 7 = 2*4 + 8=

OpenStudy (xmoses1):

That will help you ^^